OpenStudy (anonymous):

Simplify the complex fraction (2-2/x)/(3-1/x)

OpenStudy (crashonce):

With this question expand the brackets then collect like terms and multiply the entire equation by x^2

OpenStudy (crashonce):

then solve the equation simple quadratic :) CO

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (crashonce):

solve the quadratic from 6x^2 - 8x - 2 thats after you expand and collect like terms

OpenStudy (crashonce):

you might need th quadratic formula for this one

OpenStudy (skullpatrol):

$$\Huge \frac{(2-\frac{2}{x})}{(3-\frac{1}{x})}$$

OpenStudy (crashonce):

thats right

OpenStudy (skullpatrol):

$$\Huge \frac{\frac{2\cdot{x}}{x}-\frac{2}{x}}{\frac{3\cdot{x}}x-\frac{1}{x}}$$

OpenStudy (crashonce):

lol @skullpatrol yours is definitely better dont know what i did wrong

OpenStudy (skullpatrol):

$$\Huge \frac{(\frac{2\cdot{x}}{x}-\frac{2}{x})\cdot{x}}{(\frac{3\cdot{x}}x-\frac{1}{x})\cdot{x}}$$

OpenStudy (skullpatrol):

$$\Huge \frac{{2{x}}-{2}}{{3{x}}-{1}}$$